A photon of wavelength 6000nm scatters from an electron at rest. The electron recoils with an energy of 60Kev . Calculate the energy of scattered photon and the angle through which it is scattered.
1
Expert's answer
2020-04-21T19:07:54-0400

Energy before scattering:


"E_{\gamma_1}=\frac{hc}{\lambda_{\gamma_1}}=3.31\cdot10^{-20}\text{ J, or 0.207 eV}."


According to law of conservation of energy:


"E_{\gamma1}=E_{\gamma2}+Te.\\
0.207=E_{\gamma2}+60\cdot10^3,\\
E_{\gamma2}=0.207-60000<0."


This is a bit impossible and, therefore, we cannot calculate the angle of scattering.

However, assume that the wavelength of the photon is 6 pm:


"E_{\gamma_1}=\frac{hc}{\lambda_{\gamma_1}e}=206.78\text{ keV}.\\
\space\\
E_{\gamma1}=E_{\gamma2}+Te.\\
206.78=E_{\gamma2}+60,\\
E_{\gamma2}=206.78-60=146.78\text{ keV}."

The wavelength after scattering will be


"\lambda_{\gamma2}=\frac{hc}{E_{\gamma2}}=8.40\text{ pm}."


The angle through which it is scattered:


"\lambda_{\gamma2}-\lambda_{\gamma1}=\frac{hc}{m_ec^2}(1-\text{cos}\theta),\\
\space\\
\theta=\text{arccos}\bigg[1-\frac{m_ec}{h}(\lambda_{\gamma2}-\lambda_{\gamma1})\bigg]=89.3^\circ."